OpenStudy (anonymous):

Ok. Now which of the following describes the intersection of x+y+z=3 and x+y+z=6? a.Intersect at a point b.at a line c.dont intersect d.they are the same plane

OpenStudy (anonymous):

So these are planes, OK? And the normal vectors are parallel, do you know what that means.?

OpenStudy (anonymous):

Nope. I have to teach myself this so i am lost about vectors but i know what a parallel is

OpenStudy (anonymous):

Well, I don't know how you would know this but these two equations are for parallel planes.

OpenStudy (anonymous):

Which mean they don't intersect then since there parallel

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

But how would you know that?

OpenStudy (anonymous):

I don't know i guess because you said they were parallel but how did you know that they were

OpenStudy (anonymous):

Because their normal vectors are the same. <1,1,1>

OpenStudy (anonymous):

normal meaning perpendicular

OpenStudy (anonymous):

For both because theirs nothing in-front of the x y or z. And the number at the end just tell you how wide it is. I am guessing

OpenStudy (anonymous):

Close enough. They are not really 3 units apart but that's another story. Main thing is they are parallel planes.

OpenStudy (anonymous):

So what does the number behind the equal sign means to do.

OpenStudy (anonymous):

I would leave that behind for a minute. What do you know about planes and how would you answer this question with no help?

OpenStudy (anonymous):

I guess you could do that two planes will never intersect at a point two planes that are the same will have the same equation (within a constant)

OpenStudy (anonymous):

I know how to find the distance if the x,y is given. But nothing really once they add a z....

OpenStudy (anonymous):

Well it seems crazy to me. My advice: go with what we have, and if you have a new Q about planes, ping me OK?

OpenStudy (anonymous):

Ok. Well my next question is.. The graph of -3x+2y=-6z=18 intersects at which three points. Now how do you solve this

OpenStudy (anonymous):

intersects what? and are you sure -3x+2y=-6z=18 ??

OpenStudy (anonymous):

be back in 10 or 15

OpenStudy (anonymous):

ok and i meant -3x+2y-6z=18

OpenStudy (mathmate):

Do you mean intersects the coordinate axes at which 3 points?

OpenStudy (mathmate):

To find where it intersects the x-axis, set y=z=0 to get -3x+2(0)-6(0)=18, so x=-6. The intersection point is therefore (-6,0,0). You can do the same for the two other axes.

OpenStudy (anonymous):

so just plug in zero for all of them and ill find my answer

OpenStudy (mathmate):

Not all of them, just two at a time and solve for the remaining variable. Please first confirm that you do want the intersection of the plane with the axes.

OpenStudy (anonymous):

ok i get it now

OpenStudy (mathmate):

Great! Is it clear how you would find out if two planes are parallel?

OpenStudy (anonymous):

nope

OpenStudy (mathmate):

You know how to find the normal vector (like the <1,1,1> that you mentioned)?

OpenStudy (anonymous):

Yes because i know that by infront of x is a 1

OpenStudy (mathmate):

Basically, that's what it is. Put the equation of the plane into the form ax+by+cz=d the normal vector would be <a,b,c>. So far so good?

OpenStudy (anonymous):

Yes thats the basic

OpenStudy (mathmate):

For two given planes, you would find the normal vectors for each one, say <a,b,c> and <p,q,r>. Check if p/a=q/b=r/c (assuming they are all non-zero). If equality holds, then the two planes are parallel. Example: Is 2x+3y+4z=6 parallel to -4x-6y-8z=-12?

OpenStudy (anonymous):

ok ok i see

OpenStudy (mathmate):

So are they parallel?

OpenStudy (anonymous):

No but i dont know how to graph it

OpenStudy (mathmate):

In fact, they are parallel, because -4/2 = -6/3 = -8/4 (=-2) So the two planes are actually parallel.

OpenStudy (anonymous):

Ok now i am back lost this is getting kind of confusing

OpenStudy (mathmate):

The easy way to see this is compare the two normal vectors, <2,3,4> and <-4, -6, -8> If you can multiply one by a constant to get the other, then the two planes are parallel. Here the constant is -2, because -2<2,3,4>=<-4,-6,-8> therefore the two planes are parallel. Is that OK?

OpenStudy (anonymous):

Oh makes sense so they have common factor

OpenStudy (mathmate):

Kind of! I would say that if the components are proportional, then they are parallel. So far so good?

OpenStudy (anonymous):

Yes thanks for slowing it down for me

OpenStudy (mathmate):

It is important that you understand every step. That's why I am here. Next: do you know how to find out if the two planes coincide? That is if the two planes have infinite points of intersection?

OpenStudy (anonymous):

Oh no whats that

OpenStudy (mathmate):

If two parallel planes are so close to each other that they actually coincide. This is a special situation where the intersection of the two planes is not a line any more, but a plane!

OpenStudy (anonymous):

so if they are real close then its a coincide but how far part of way

OpenStudy (mathmate):

To check if they are coincident, you would multiply one of the equations by the factor k to make the left-hand sides equal. If the right-hand sides also equal, then the two planes coincide, or are identical. Take the same two planes we had: 2x+3y+4z=6 ...(1) -4x-6y-8z=-12 ....(2) Multiplying plane (1) by -2, we get -4x-6y-8z=-12 ...(3) The left hand side of (3) is the same as plane 2, so (1) and (2) are parallel (as before). The right-hand side of -12 is also equal to the right-hand side of (2), so (1) and (2) are coincident, and the intersection is a plane, namely plane (1) or plane (2).

OpenStudy (anonymous):

ooooooooooooooo